# Towards Scalable IoT Deployment for Visual Anomaly Detection via Efficient Compression

### v2

@@ 1 Introduction
VAD research ranges from computer vision to deep learning, as documented in surveys [1]. Studies have explored anomaly detection under resource constraints [5]. Edge-assisted systems reduce latency in industrial monitoring [2], while [3] proposes low-latency VAD for smart cities. Research has also explored edge computing integration [4], traffic flow analysis [6] and health monitoring systems [7].
@@
While a universally accepted taxonomy of VAD methods is lacking, recent studies [14, 15] propose a classification into three main categories: reconstruction-based, synthesizing-based, and embedding-based approaches. Reconstruction-based methods assume anomalies cannot be reconstructed due to their absence in the training data (e.g., visual sensor networks in [16]; they use autoencoders, GANs, diffusion models, and Transformers to reconstruct normal samples, identifying anomalies via pixel-wise or SSIM-based differences. However, overfitting or visually similar anomalies can lead to missed detections. Synthesizing-based methods generate artificial anomalies on normal images, such as CutPaste [17] and DRÆM [18], but struggle to capture real-world defect variability. Finally, feature-based methods use pre-trained models to represent normal data distributions; examples include PaDiM [19], PatchCore [20], and FastFlow [21]. These methods achieve high accuracy but can be resource-intensive and sensitive to domain shifts.
Nonetheless, to the best of our knowledge, the existing literature lacks solutions that jointly optimize VAD models and communication strategies for IoT. This paper proposes a lightweight, feature-based VAD framework for bandwidth-constrained IIoT environments, emphasizing efficient data transmission and interpretable anomaly localization. Our contributions are as follows:
  • We propose a modular, resource-aware VAD pipeline for distributed edge-server deployment, optimizing detection performance, model footprint, and communication overhead for efficient anomaly detection in IoT environments.
  • We investigate two strategies: edge image compression and lightweight CNN feature extraction with compression, reducing computational and communication burdens while preserving detection accuracy. Also, we explore compression techniques like WebP for image encoding, and random projection and product quantization for feature encoding, achieving high compression rates with minimal performance loss.
@@ 4.2 Examined scenarios
All configurations are based on PatchCore [20] as the anomaly detection method, and all input images from the MVTec AD dataset [24] are uniformly resized to $224\times 224$. This resolution is standard in the VAD literature e.g., PaDiM [19], CFA [25] and is compatible with the input sizes of the MobileNetV2 and WideResNet backbones used in our experiments.
@@
  • Raw Image Transmission ($224\times 224$): The downsampled image is sent directly to the server, where WideResNet performs feature extraction and anomaly detection.
@@ 5 Conclusions and future directions
We have proposed a modular and efficient visual anomaly detection (VAD) framework tailored for IoT scenarios, where edge resource constraints pose major challenges. By investigating both image and feature compression strategies, we show that communication overhead can be significantly reduced while maintaining high detection accuracy. Validated on the MVTec AD dataset using a lightweight CNN and PatchCore, our approach matches the performance of uncompressed pipelines with lower latency and memory usage, making it viable for industrial deployment.
Future work will explore dynamic adaptation strategies that adjust compression based on real-time factors like bandwidth and computational load. We also plan to investigate learning-based feature compression and extend the framework to support online model updates and real-time anomaly localization across heterogeneous edge-server systems.

### v3

@@ 1 Introduction
VAD research ranges from computer vision to deep learning, as documented in surveys [1]. Studies have explored anomaly detection under resource constraints [5]. Edge-assisted systems reduce latency in industrial monitoring [2], while [3] proposes low-latency VAD for smart cities. Research has also explored edge computing integration [4], traffic flow analysis [6], health monitoring systems [7] and data-driven predictive control [8].
@@
While a standard taxonomy of VAD methods is lacking, recent studies [15, 16] propose a classification into three categories: reconstruction-, synthesizing-, and embedding-based approaches. Reconstruction methods assume anomalies are poorly reconstructed due to their absence in the training data (e.g., visual sensor networks in [17]); using autoencoders, GANs, diffusion models, and Transformers to reconstruct samples, identifying anomalies via pixel-wise or SSIM differences. However, overfitting or visually similar anomalies can lead methods to underperform. Synthesizing methods generate artificial anomalies on normal images, such as CutPaste [18] and DRÆM [19], but sometimes struggle to capture real-world defect characteristics. Finally, embedding methods use pretrained models to represent normal data distributions; e.g. PaDiM [20], PatchCore [21], FastFlow [22]. These methods achieve high accuracy but can be resource-intensive and sensitive to domain shifts.
Nonetheless, to the best of our knowledge, existing literature lacks solutions that jointly optimize VAD models and communication strategies for IoT. This paper proposes a lightweight, embedding-based VAD framework for bandwidth-constrained IIoT environments, with efficient data transmission and interpretable anomaly localization. Our contributions are as follows:
  • Resource-aware VAD pipeline for distributed edge-server deployment, optimizing detection performance, model footprint, and communication overhead for efficient anomaly detection in IoT environments.
  • Test two strategies: edge image compression VS tiny CNN feature extraction with compression, reducing computational and communication resources while preserving VAD performance. We test compression methods like WebP for images, random projection and product quantization for feature encoding.
@@ 4.2 Examined scenarios
All configurations are based on PatchCore [21] as the anomaly detection method, and all input images from the MVTec AD dataset [25] are uniformly resized to $\times 224$ 224. This resolution is standard in the VAD literature (e.g., PaDiM ([20]), CFA [26]) and is compatible with the input sizes of the MobileNetV2 and WideResNet backbones used in our experiments.
@@
  • Raw Image Transmission ($\times 224$ 224): The downsampled image is sent directly to the server, where WideResNet performs feature extraction and anomaly detection.
@@ 5 Conclusions and future directions
We have proposed an efficient VAD framework for IoT scenarios, included in our open-source library MoViAD [^1] https://github.com/AMCO-UniPD/moviad [27]. Testing both image and feature compression strategies shows that communication overhead can be reduced while maintaining high detection accuracy. Validated on the MVTec AD dataset using a lightweight CNN and PatchCore, our approach matches the performance of uncompressed pipelines with lower latency and memory usage, making it viable for industrial deployment. Future work will explore dynamic adaptation strategies that adjust compression based on real-time factors like bandwidth and computational load. We also plan to investigate learning-based feature compression and extend the framework to support online model updates and real-time anomaly localization across heterogeneous edge-server systems. Another promising direction involves embedding security mechanisms directly into the framework, aiming to enhance the resilience of heterogeneous edge-server networks [28, 29].